What is the density of unknown gas with a specific gravity of 0.8 at STP?


1
Expert's answer
2020-09-21T06:56:51-0400

According to the equation:

SG = d(gas) / d(water)

where SG - specific gravity, d - density.

From here, the density of unknown gas equals:

d(gas) = SG × d(water)

As the density of water equals 1 g/ml, the density of unknown gas is:

d(gas) = 0.8 × 1 g/ml = 0.8 g/ml


Answer: 0.8 g/ml